Gasoline : 1 , 2 , 3 , 4 , 5

(a) Unleaded : $2.27, $2.27, $2.27, $2.27, $2.27

(b) Unleaded plus : 2.36, 2.37, 2.38, 2.38, 2.40

(c) Super Unleaded: 2.47, 2.49, 2.50, 2.50, 2.59


i) What is the mean price of: (a), (b), and (c).                                        

 

ii) What is the median price of: (a), and (b).

                                         

iii) What is the mode price of: (a), and (b)


1
Expert's answer
2021-07-13T13:01:09-0400

(i)

(a)

"mean=\dfrac{2.27+2.27+2.27+2.27+2.27}{5}=\$2.27"

(b)

"mean=\dfrac{2.36+2.37+2.38+2.38+2.40}{5}=\$2.38"

(c)

"mean=\dfrac{2.47+2.49+2.50+2.50+2.59}{5}=\$2.51"

(ii)

(a)

 $2.27, $2.27, $2.27, $2.27, $2.27

"median=\$2.27"

(b)

$2.36, $2.37, $2.38, $2.38, $2.40

"median=\$2.38"

(c)

 $2.47, $2.49, $2.50, $2.50, $2.59

"median=\$2.50"

(iii)

(a)  $2.27 occurs 5 times

"mode=\$2.27"

(b)

$2.36 occurs 1 time

$2.37 occurs 1 time

$2.38 occurs 2 times

$2.40 occurs 1 time

"mode=\$2.38"

(c)

$2.47 occurs 1 time

$2.49 occurs 1 time

$2.50 occurs 2 times

$2.59 occurs 1 time

"mode=\$2.50"